# Category $\mathcal{O}$ for hybrid quantum groups and non-commutative Springer resolutions

### v3

@@ 7.2.3. Graded multiplicities
Since $\widetilde{{\mathcal{Q}}}^{\leq\nu}_{x}$ is the truncation of $\widetilde{{\mathcal{Q}}}_{x}$ by $\nu$, (7.7) follows.
Now we study the multiplicity [~$[\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}:\widetilde{{\mathcal{L}}}_{y}]_{\mathrm{gr}}$. Consider the following sequence induced by Koszul resolution
(7.11) ~$\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}\otimes_{{\mathbb{C}}[{\mathfrak{n}}]}({\mathbb{C}}[{\mathfrak{n}}]\otimes\bigwedge^{\bullet}{\mathfrak{n}}^{*})\rightarrow\widetilde{{\mathcal{Z}}}_{x}\rightarrow 0,$
is exact by the freeness of the $Z_{\mathrm{Fr}}^{-}$-module $M(x\bullet_{l}0)_{\mathbb{C}}$. By abuse of notation, we view $\langle\eta\rangle$ ($\eta\in\Lambda$) as a continuous ${\mathbb{Z}}[v^{\pm 1}]$-linear operator on $\hat{K}_{0}({\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}})$ sending [~$[\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}]$ to [~$[\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{t(\eta)x}]$, for any $x\in W_{\mathrm{ex}}$. By (7.11), we have the following equality in $K_{0}({\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}})$,
(7.12) $[\widetilde{{\mathcal{Z}}}_{x}]=\big{(}\prod_{\alpha\in\Phi^{+}}(1-v^{2}\langle-\alpha\rangle)\big{)}[\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}].$
In $\hat{K}_{0}({\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}})$, we express [~$[\widetilde{{\mathcal{L}}}_{y}]$ into (topological) basis $\{[\widetilde{{\mathcal{Z}}}_{x}]\}_{x}$ and {[~$\{[\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}]\}_{x}$ as
[~$[\widetilde{{\mathcal{L}}}_{y}]=\sum_{x}[\widetilde{{\mathcal{L}}}_{y}:\widetilde{{\mathcal{Z}}}_{x}]_{\mathrm{gr}}\cdot[\widetilde{{\mathcal{Z}}}_{x}]=\sum_{x}[\widetilde{{\mathcal{L}}}_{y}:\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}]_{\mathrm{gr}}\cdot[\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}].$
Then (7.12) shows that
By [29, §11] (see also [38, Thm 6.1]), there is an equality
(7.14) $\sum_{x}(-1)^{l(x)+l(y)}q_{x,y}(v)\cdot p_{w_{0}x,w_{0}z}(v)=\delta_{y,z},\quad\forall y,z\in W_{\mathrm{ex}}.$
Hence it follows from Proposition 7.6 that
(7.15) [~$[\widetilde{{\mathcal{L}}}_{y}:\widetilde{{\mathcal{Z}}}_{x}]_{\mathrm{gr}}=(-1)^{l(x)+l(y)}q_{x,y}(v),\quad\forall x,y\in W_{\mathrm{ex}}.$
Combining (7.1), (7.13) with (7.15), we deduce that
[~$[\widetilde{{\mathcal{L}}}_{y}:\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}]_{\mathrm{gr}}=(-1)^{l(x)+l(y)}p_{x,y}(v),\quad\forall x,y\in W_{\mathrm{ex}}.$
By (7.14) again, we have
[~$[\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}:\widetilde{{\mathcal{L}}}_{y}]_{\mathrm{gr}}=q_{w_{0}x,w_{0}y}(v),\quad\forall x,y\in W_{\mathrm{ex}}.\qed$

### v4

@@ 7.2.3. Graded multiplicities
To complete the proof, we show that as a truncation of $\widetilde{{\mathcal{Q}}}_{x}$, the module $\widetilde{{\mathcal{Q}}}^{\Omega}_{x}$ admits graded multiplicity obtained as truncation from the one for $\widetilde{{\mathcal{Q}}}_{x}$, as (7.9). As $(T\times{\mathbb{C}}^{\times})$-equivariant ${\mathsf{A}}_{\mathfrak{n}}$-modules, we have $\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{y}\otimes U{\mathfrak{n}}\simeq\bigoplus_{\eta}\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{y}\otimes(U{\mathfrak{n}})_{\eta}$. Since $\mathrm{Ext}^{1}(\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x},\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{y}\langle d\rangle)=0$ if $x\bullet_{l}0\nless y\bullet_{l}0$, from the filtration (7.19) one can obtain a separated decreasing filtration of $\widetilde{{\mathcal{Q}}}_{x}$ in ${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}^{T\times{\mathbb{C}}^{\times}}$, with composition factors by Verma objects. Hence $\tau^{T\times{\mathbb{C}}^{\times},\Omega}(\widetilde{{\mathcal{Q}}}_{x})$ (see again the proof of Lemma 7.3 for the truncation functor $\tau^{T\times{\mathbb{C}}^{\times},\Omega}$) is the quotient of $\widetilde{{\mathcal{Q}}}_{x}$ by the submodule composed by $\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{y}\langle d\rangle$ with $y\bullet_{l}0\notin\Omega$, and therefore its graded multiplicity of Verma objects is as the RHS of (7.9). Since $\widetilde{{\mathcal{Q}}}_{x}^{\Omega}=\tau^{U{\mathfrak{n}}\times{\mathbb{C}}^{\times},\Omega}(\widetilde{{\mathcal{Q}}}_{x})$ is a quotient of $\tau^{T\times{\mathbb{C}}^{\times},\Omega}(\widetilde{{\mathcal{Q}}}_{x})$, we have
[…]
On the other hand, by the BGG reciprocity [35, Prop 3.4] (in loc. cit. it is stated for $\Omega=\{\lambda\in\Lambda|\lambda\leq\nu\}$ for some $\nu$, but the proof works for general $\Omega$) and (7.17) we have
[…]
Hence (7.20) is an equality. It follows that $\widetilde{{\mathcal{Q}}}_{x}^{\Omega}=\tau^{U{\mathfrak{n}}\times{\mathbb{C}}^{\times},\Omega}(\widetilde{{\mathcal{Q}}}_{x})=\tau^{T\times{\mathbb{C}}^{\times},\Omega}(\widetilde{{\mathcal{Q}}}_{x})$, whose graded multiplicity $(\widetilde{{\mathcal{Q}}}_{x}^{\Omega}:\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{y})_{\mathrm{gr}}=(\tau^{T\times{\mathbb{C}}^{\times},\Omega}(\widetilde{{\mathcal{Q}}}_{x}):\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{y})_{\mathrm{gr}}$ is equal to the RHS of (7.9). ∎